OpenStudy (anonymous):

Find the probability that a 5 person team consists of 3 boys and 2 girls if the team is chosen at random from 5 girls and 3 boys.

OpenStudy (anonymous):

The number of combinations that give three girls and two guys is 5C3 times 3C2. The total number of combinations is 8C5, so the probability is\[\frac{\left(\begin{matrix}5 \\ 3\end{matrix}\right)\left(\begin{matrix}3 \\2\end{matrix}\right)}{\left(\begin{matrix}8 \\ 5\end{matrix}\right)}\]

OpenStudy (anonymous):

I think it's about 30/56=15/28

OpenStudy (anonymous):

Got it?